OpenStudy (anonymous):

what is the equation of a line that passes through the point (0,5) and has a negative slope

OpenStudy (danjs):

point slope: (y - y1) = m* (x - x1)

OpenStudy (danjs):

where (x1,y1) = (0,5)

OpenStudy (danjs):

y - 5 = m*(x-0) y = m*x +5

OpenStudy (danjs):

m is any negative value

OpenStudy (anonymous):

thank you